# Matrix-Driven Instant Review: Confident Detection and Reconstruction of LLM Plagiarism on PC

### v2

@@ 3.6 Estimating the $p$-Value
For the proof, please refer to Appendix A.
Based on this result, our $p$-value can now be estimated as follows:
$p$ $\leq n!\cdot f(c)$
$\approx n!\cdot P(c/n,m)$
$\asymp n!\cdot\exp(-2m^{2}c^{2}/n^{2})$
$=n!\cdot\exp\left(-\frac{c^{2}}{2}\right).$
However, this estimation remains coarse and imprecise. To provide definitive statistical evidence of model similarity, we recommend adopting a threshold of at least ten standard deviations ($10\sigma$), corresponding to a $p$-value of less than $2\times 10^{-23}$. Achieving this criterion is feasible in the context of billion-parameter-scale models, as even tiniest similarities will yield extremely significant results.
Refining the $p$-value estimation further would require advanced mathematical tools and techniques, which are beyond the scope of this research. This topic may warrant a separate investigation in future work.
3.7 Solving the Rest
If the significance of the $p$-value has already been determined from the previous calculations, this step is entirely optional. However, if the $p$-value is not significant at this stage, we cannot yet rule out the possibility of plagiarism. This may occur when the outer transformation is a general orthogonal matrix. Additionally, interested readers may wish to determine the exact relationship between two models, potentially with different architectures.
Since $P=\arg\max\operatorname{Tr}(P\tilde{U}^{T})$, we proceed as follows: If $P$ is reliably identified as a permutation matrix ($p<2\times 10^{-23}$), set $U:=P$; otherwise, set $U:=\tilde{U}$.
Our objectives are:
$\min_{W_{Q}\text{ orthogonal}}\lVert UQW_{Q}-Q^{\prime}\rVert_{F}^{2},\quad\min_{W_{K}\text{ orthogonal}}\lVert UKW_{K}-K^{\prime}\rVert_{F}^{2},$
$W_{Q}=\lambda\mathrm{Ortho}(Q^{T}U^{T}Q^{\prime}),\quad W_{K}=\mu\mathrm{Ortho}(K^{T}U^{T}K^{\prime}),$
where the scaling coefficients $\lambda$ and $\mu$ are computed as:
For $W_{V}$ and $W_{O}$, if the model stealer uses an orthogonal transformation, we apply the same method to compute $\mathrm{Ortho}(V^{T}U^{T}V^{\prime})$ and $\mathrm{Ortho}(OU^{T}O^{\prime T})$. However, since $W_{V}$ and $W_{O}$ may involve general invertible transformations, solving the general case without additional assumptions is challenging. We leave this problem for future work.
$U_{X}$ $=\mathrm{Ortho}(X^{T}U^{T}X^{\prime}),\quad X\in\{\mathrm{Gate},\mathrm{Up},\mathrm{Down}\};$
$P$ $=\arg\max_{P\in\mathrm{Perm}(\mathrm{IntermediateDim})}\operatorname{Tr}\left(P\left(U_{\mathrm{Gate}}+U_{\mathrm{Up}}+U_{\mathrm{Down}}\right)^{T}\right).$
Typically, we expect the three solutions $\arg\max\operatorname{Tr}\left(PU_{X}^{T}\right)$, for $X\in\{\mathrm{Gate},\mathrm{Up},\mathrm{Down}\}$, to yield the same permutation. However, computing the orthonormal part for intermediate matrices (which often have 14,000–20,000 rows) is computationally expensive. Adding the three terms together would triple the computation cost. If the noise level is tolerable, we may simply select one of them:
$P=\arg\max_{P\in\mathrm{Perm}(\mathrm{IntermediateDim})}\operatorname{Tr}\left(PU_{\mathrm{Up}}^{T}\right).$
Similarly, the $p$-value for permuted MLPs can be estimated as:
$x$ $=\operatorname{Tr}\left(PU_{\mathrm{Up}}^{T}\right),$
$p_{\mathrm{MLP}}$ $=\mathrm{IntermediateDim}!\cdot\exp(-x^{2}/2).$
However, this is not an accurate estimation for the MLP, as similar features may gradually emerge during training, and the correspondence between neurons may also arise independently.
3.8 Overall Algorithm
We summarize our algorithm as follows:
Algorithm 1 Computing the $p$-value: $\mathrm{PValue}(P,U,d)$
Input: Permutation matrix $P$, orthogonal matrix $U$, dimension $d$
Compute $p:=d!\cdot\exp(-(\operatorname{Tr}(PU^{T}))^{2}/2)$
Output: $p$
Algorithm 2 MDIR (Matrix-Driven Instant Review)
Input: Model $A$, Model $B$
Initialize: Threshold $p_{0}=2\times 10^{-23}$, relation flag $r=\mathsf{False}$
$L:=$ Number of layers in the model
$E:=$ Embedding matrix of AA
$E^{\prime}:=$ Embedding matrix of $B$
$\mathcal{C}:=$ Set of common vocabulary tokens between $A$ and $B$
Compute $\tilde{U}:=\mathrm{Ortho}(E[\mathcal{C},:]^{T}E^{\prime}[\mathcal{C},:])$ via polar decomposition
$P:=\arg\max_{P\in\mathrm{Perm}(\mathrm{EmbDim})}\operatorname{Tr}(P\tilde{U}^{T})$
Compute $p:=\mathrm{PValue}(P,\tilde{U},\mathrm{EmbDim})$
if $p<p_{0}$ then
  Set $r:=\mathsf{True}$
  Set $U:=P$
else
  Set $U:=\tilde{U}$
end if
Yield: Transformation matrix $U$, $p$-value $p$
for Layer $i\in[L]$ do
  Extract attention weights $Q,K,V,O$ from layer $i$ of AA
  Extract attention weights $Q^{\prime},K^{\prime},V^{\prime},O^{\prime}$ from layer $i$ of $B$
  for $X\in\{Q,K,V\}$ do
   Compute $W_{X}:=\mathrm{Ortho}(X^{T}U^{T}X^{\prime})\cdot\frac{\lVert X^{\prime}\rVert_{F}}{\lVert X\rVert_{F}}$
  end for
  Extract MLP weights $\mathrm{Gate},\mathrm{Up},\mathrm{Down}$ from layer $i$ of AA
  Extract MLP weights $\mathrm{Gate}^{\prime},\mathrm{Up}^{\prime},\mathrm{Down}^{\prime}$ from layer $i$ of $B$
  Compute $U_{X}:=\mathrm{Ortho}(X^{T}U^{T}X^{\prime})$ for $X\in\{\mathrm{Up}\}$ (or $X\in\{\mathrm{Gate},\mathrm{Up},\mathrm{Down}\}$)
  $P:=\arg\max_{P\in\mathrm{Perm}(\mathrm{IntermediateDim})}\operatorname{Tr}(PU_{\mathrm{Up}}^{T})$
  Compute $p:=L\cdot\mathrm{PValue}(P,U_{\mathrm{Up}},\mathrm{IntermediateDim})$
  if $p<p_{0}$ then
   Yield: Transformation matrix $U_{\mathrm{Up}}$, $p$-value $p$
   Set $r:=\mathsf{True}$
  end if
end for
Output: Relation flag $r$
4 Experiments and Case Studies
We conducted a series of experiments to evaluate the effectiveness of our MDIR method in identifying similarities between models. Considering file size constraints, for each matrix, we plotted the first $512\times 512$ submatrix. The complete plots of the entire matrices can be produced by the accompanying code.
4.1 Level 1: Identifying Instruction-tuned Versions and Avoiding False Positives for Unrelated Models
This section examines official instruction-tuned versions of models in comparison to their base counterparts:
  • Qwen2.5-0.5B and Qwen2.5-0.5B-Instruct [34]: Figure LABEL:fig:qwen25_0b5
  • Meta-Llama-3.1-8B and Meta-Llama-3.1-8B-Instruct [28]: Figure LABEL:fig:llama31_8b
These pairs of models are known to be unrelated, even through they share similar architectures:
  • Meta-Llama-3.1-8B and Qwen3-8B-Base [51]: Figure LABEL:fig:llama3_qwen3
  • DeepSeek-V3-Base [10] and Kimi-K2-Instruct [30]: Figure LABEL:fig:ds_kimi
4.2 Level 2: Identifying Upcycling Models
These models are known as upcycled from their predecessors, according to their technical reports:
  • Qwen1.5-1.8B [44] and Qwen1.5-MoE-A2.7B [45]: Figure LABEL:fig:qwen15_18
  • Qwen2-7B and Qwen2-57B-A14B [50]: Figure LABEL:fig:qwen2_57
4.3 Level 3: Identifying Continual Pretraining of Pruned and Distilled Models
These models are pruned versions of their predecessors and have subsequently undergone continual pretraining:
  • Meta-Llama-3.1-8B and Meta-Llama-3.2-1B [1]: Figure LABEL:fig:llama31_llama32_1b
  • Meta-Llama-3.1-8B and Meta-Llama-3.2-3B: Figure LABEL:fig:llama31_llama32_3b
We observe that the large values, shown by red pixels, form irregular sloped lines in the relation matrices. Each row (from the smaller model) corresponds to a single large value associated with one channel from the larger model. This pattern indicates that the smaller model is a pruned version of the larger model.
Our computed $p$-value is heavily overestimated, and the actual $p$-value should be orders of magnitude smaller. Since the channels of the smaller model are arranged in an orderly manner, there are only $\mathbf{C}(n,m)$ admissible configurations for such a pattern, not $n!$. Unfortunately, we have not derived the correct large deviation principle for non-square semi-orthogonal matrices like in this example. Therefore, we decide to retain this over-estimated $p$-value, and leave the refinement of the analysis to future work.
4.4 Level 4: Identifying Continual Pretraining of Non-Transformer Models with Architectural Modifications
In addition to Transformer-based models, our work also detects relationships between models of non-Transformer architectures:
  • RWKV-7-World-0.4B [32] and RWKV-5-Eagle-0.4B [31]: Figure LABEL:fig:rwkv57
  • RWKV-7-World-2.9B and RWKV-6-Finch-3B: Figure LABEL:fig:rwkv67
The RWKV series (RWKV-5, 6 and 7) uses a head size of $64\times 64$, which differs from a traditional $128\times 128$ head size commonly used in most modern Transformers. This $64\times 64$ block pattern is clearly visible from Figure LABEL:fig:rwkv57(b). However, with significant architectural changes, the diagonal patterns become less distinct and more challenging to identify.
4.5 Level 5: Identifying Plagiarism in Continual Pretraining of Upscaled Models with Comprehensive Obfuscation
A recent debate has emerged regarding whether the Pangu-Pro-MoE model [6, 5] is derived from Qwen2.5-14B. Our analysis provides strong evidence supporting this claim with a high level of confidence.
  • Pangu-Pro-MoE [6] and Qwen2.5-14B: Figure LABEL:fig:pangu_qwen
To solve the relationship between the embedding channels, we identified 101,486 tokens in common between the tokenizer of Pangu-Pro-MoE (which contains 153,376 tokens) and that of Qwen2.5-14B (with 151,665 tokens). The initial step in solving the relation matrix from these vocabulary embeddings of these overlapping tokens yielded a trace of 3862, corresponding to a $p$-value less than $10^{-3,000,000}$. Subsequent analyses consistently produced similarly low $p$-values, indicating a significant similarity in the embeddings, attention modules, and shared expert/MLP modules between the two models. As illustrated in Figure LABEL:fig:pangu_qwen, the red dots in the nearly white background, though barely visible, reveal that these matrices are nearly permutation matrices. Taking into account the temporal and contextual factors, such as the fact that Pangu-Pro-MoE was released approximately nine months after Qwen2.5-14B, we conclude that Pangu-Pro-MoE reuses weights from Qwen2.5-14B. This finding raises concerns about the validity and reliability of the technical report of Pangu-Pro-MoE [5].
4.6 Ablation: MDIR Only Detects Weight Relevance, Not Training Data
We demonstrate that our MDIR method exclusively detects relevance in weights, rather than training data. To validate this claim, we conducted an ablation experiment by initializing two models with different random seeds and training each on two distinct datasets, resulting in a total of four models.
The datasets used were:
  • DCLM subsample [24] (the first 100 files of shard 0 [29], containing 12.05 billion tokens), and
  • OpenWebMath-ProX [56] (4.61 billion tokens [15], abbreviated as OWM hereafter).
Both datasets were tokenized using the GPT-NeoX tokenizer [7].
[…]
Figure LABEL:fig:seed reveals clear diagonal patterns for models initialized with the same seed, indicating strong weight similarity due to shared initialization. In contrast, Figure LABEL:fig:samedata shows varying brightness blocks for solved attention relation matrices but no significant outliers, suggesting that models trained on the same dataset may develop similar attention features without substantial weight correlation.
4.7 Comparative Analysis
REEF Method May Produce False Positives. We compare our method against REEF (Representation Encoding Fingerprint) [55] using the following models:
  • Meta-Llama-3.1-8B vs. Meta-Llama-3.2-3B: True positive (Figure LABEL:fig:llama31_llama32_3b),
  • Meta-Llama-3.1-8B vs. Qwen2.5-7B: False positive (Figure LABEL:fig:reef).
The results indicate a similarity score exceeding $0.8$ for the first few layers (e.g., layer 2 of Meta-Llama-3.1-8B vs. layer 2 of Qwen2.5-7B), despite the fact that the Llama 3.1 and Qwen2.5 series are entirely unrelated.
This phenomenon may arise because both models, after training on tens of trillions of tokens, gradually develop similar representations, which can be detected by CKA (Central Kernel Alignment). In contrast, our MDIR method is specifically designed to detect weight similarity, not representation similarity, and safely returns negative results in such cases (Figure LABEL:fig:comparison).
5 Conclusion
We introduced MDIR, a novel method for LLM weight similarity and plagiarism detection, leveraging robust mathematical tools and providing statistically meaningful results. Our method does not require full model inference and can be executed on a single PC or even without a GPU, democratizing the verification process. Moreover, our method detects only weight similarity but not feature similarity, making it more specific and targeted than previous techniques. This reduces the likelihood of false positives and enhances reliability.
5.1 Limitations
Numerical Precision. All of our discussions above assume infinite precision; however, practically, our method relies on matrix polar decomposition, which is an ill-posed problem when the matrix is low-rank. Moreover, since most neural networks are trained using low-precision formats (fp16, bf16, or even fp8), this challenges our assumption that transformation matrices are uniformly distributed across the Haar measure. In practice, we also observed approximately a 1% difference when computing the polar part under fp32 and fp64.
Extreme $p$-values. Our method produces $p$-values on the order of $10^{-3,000,000}$, far exceeding $p$-values observable in traditional statistical settings. This is reasonable given the scale of modern models, which often contain billions of parameters, compared to the hundreds or thousands of samples typically analyzed in conventional statistics. However, this may raise concerns when combined with numerical precision issues. For example, assuming a typical error of 2%, and if unfortunately $25\sigma$ of error deviates from the infinite-precision ground truth, it could literally halve the trace and diminish the $p$-value significance obtained from the trace. Therefore, the actual $p$-value, though still highly significant statistically, is much larger than that reported by our method.
5.2 Future Work
Potential Ways of Evading Detection. Although not observed in our experiments, there might be ways to evade our detection. We hypothesize that MDIR could be evaded through additional training with larger learning rates, especially for vocabulary embeddings. We leave this for future work.
More Accurate Estimation of $p$-values. We addressed the $p$-value for the orthogonal case (square matrices) using large deviation theory. However, accurate estimation of $p$-values for the semi-orthogonal case (non-square matrices, also known as partial isometries and Stiefel manifolds) remains an open challenge. This is more common because the embedding and intermediate dimensions often differ between two arbitrary models. We also leave this important refinement for future work.
6 Ethics Statement
In developing our MDIR framework, we deeply respect intellectual property rights and adhere to ethical AI practices. We strongly discourage any form of plagiarism or unauthorized fabrication of large language models. Such actions not only undermine innovation but also erode trust within the AI community.
Our work aims to promote the democratization, transparency, and responsibility of AI technologies. Therefore, we firmly oppose the misuse of our framework to conceal plagiarism in LLMs.
Appendix A Background and Proof of Theorem 1
Before we delve into the proof, it is worth discussing the challenges we encountered during the process.
We are dealing with the Circular Real Ensemble (CRE, named after [39]), which has a formulation distinct from traditional circular ensembles, such as Circular Unitary Ensembles (CUE) or Circular Orthogonal Ensembles (COE) [13]. These ensembles play important roles in both random matrix theory and condensed matter physics.
One of the problems most similar to ours is the Gross-Witten ensemble [18, 16], which concerns the large deviations from the typical value of $\mathrm{Re}\operatorname{Tr}(U)$ as $N\to\infty$ (where $U\in\mathrm{U}(N)$ is a random unitary matrix).
@@ Appendix A Background and Proof of Theorem 1
The circular ensemble is closely related to thermodynamics and large deviation theory. We refer to Mehta’s book Random Matrices [27], and a comprehensive introduction can be found in Chapter 12 of it.
A.1 Proof of Theorem 1
@@
Denote by $\{e^{i\theta_{k}},e^{-i\theta_{k}}:k=1,2,\cdots,m\}$ the eigenvalues of $A$. It is a classical result (see [48, 17]) that the phases $(\theta_{k})_{k}$ obey the distribution characterized by the following probability density:
@@
Inspired by Chapter 12 of [27] and also [46], we define the Canonical Ensemble of these $m$ particles as follows, which admits a partition function:
@@
We now solve the exact form of $F(\mu)$. We refer to Theorem 2.1 of the paper [14], where we are dealing with the case $\theta=1$, $\kappa=1$, and $w_{m}(x)=(1-x^{2})^{1/(2m)}$, where $w_{m}(x)\to 1$ in the thermodynamic limit $m\to\infty$. This suggests that the interaction term dominates, and the external field term is negligible when $m$ is large.
@@
From Corollary 2.2 of the paper [14], it suffices to solve the following two variational problems:
@@
It is known that $\log|x-y|$ has the Chebyshev expansion [42, 26]:
@@
$\quad\quad\quad\quad\left(\sum_{i=1}^{\infty}a_{i}T_{i}(x)\right)\left(\sum_{i=1}^{\infty}a_{i}T_{i}(y)\right)\mathrm{d}x\mathrm{d}y.$
@@
$F(\mu)=\log 2+\sum_{i=1}^{\infty}\frac{a_{i}^{2}}{2i}.$
@@
A.2 Additional Notes on the $\mathrm{O}(2m)\setminus\mathrm{SO}(2m)$ Case
@@
A.3 The Exact Form of $I(r)$
@@
Like the Gross-Witten Ensemble [18], there is a third-order phase transition near $r=1/2$.

### v3

@@ Appendix H Estimation of $p$-value via Large Deviation Theory
Before we delve into the proof, it is worth discussing the challenges we encountered. We are dealing with the Circular Real Ensemble (CRE, named after PhysRevLett.104.147001), which has a formulation distinct from traditional circular ensembles, such as Circular Unitary Ensembles (CUE) or Circular Orthogonal Ensembles (COE) (Dyson1962). These ensembles play important roles in both random matrix theory and condensed matter physics.
One of the problems most similar to ours is the Gross-Witten Ensemble (GrossWitten1980; gamboa2017sumruleslargedeviations), which concerns the large deviations from the typical value of $\mathrm{Re}\operatorname{Tr}(U)$ as $N\to\infty$ (where $U\in\mathrm{U}(N)$ is a random unitary matrix). Our problem can be regarded as a real orthogonal variant of Gross-Witten Ensemble, but this is undocumented in previous literature.
@@
The circular ensemble is closely related to thermodynamics and large deviation theory. We refer to Mehta’s book Random Matrices (Mehta2004), and a comprehensive introduction can be found in Chapter 12 of it.
H.1 Proof of Theorem 3
@@
Denote by $\{e^{i\theta_{k}},e^{-i\theta_{k}}:k=1,2,\cdots,m\}$ the eigenvalues of $A$. It is a classical result (see weyl2016classical; Girko1985) that the phases $(\theta_{k})_{k}$ obey the distribution characterized by the following probability density:
@@
Inspired by Chapter 12 of Mehta2004 and also TOUCHETTE20091, we define the Canonical Ensemble of these $m$ particles as follows, which admits a partition function:
@@
We now solve the exact form of $F(\mu)$. We refer to Theorem 2.1 of the paper (Eichelsbacher_2011), where we are dealing with the case $\theta=1$, $\kappa=1$, and $w_{m}(x)=(1-x^{2})^{1/(2m)}$, where $w_{m}(x)\to 1$ in the thermodynamic limit $m\to\infty$. This suggests that the interaction term dominates, and the external field term is negligible when $m$ is large.
@@
From Corollary 2.2 of the paper (Eichelsbacher_2011), it suffices to solve the following two variational problems:
@@
It is known that $\log|x-y|$ has the Chebyshev expansion (SloanStephan1992; MasonHandscomb2002):
@@
$\quad\quad\quad\quad\left(\sum_{i=1}^{\infty}a_{i}T_{i}(x)\right)\left(\sum_{i=1}^{\infty}a_{i}T_{i}(y)\right)\mathrm{d}x\mathrm{d}y$
$\quad+\iint_{[-1,1]^{2}}\left(\sum_{n=1}^{\infty}\frac{2}{n}\frac{T_{n}(x)\left(\sum_{i=1}^{\infty}a_{i}T_{i}(x)\right)}{\pi\sqrt{1-x^{2}}}\frac{T_{n}(y)\left(\sum_{i=1}^{\infty}a_{i}T_{i}(y)\right)}{\pi\sqrt{1-y^{2}}}\right)\mathrm{d}x\mathrm{d}y$
$=\log 2\left(\int_{-1}^{1}\frac{1}{\pi\sqrt{1-x^{2}}}\mathrm{d}x\right)^{2}$
$\quad+\sum_{n=1}^{\infty}\frac{2}{n}\left(\int_{-1}^{1}\frac{T_{n}(x)\left(\sum_{i=1}^{\infty}a_{i}T_{i}(x)\right)}{\pi\sqrt{1-x^{2}}}\mathrm{d}x\right)\left(\int_{-1}^{1}\frac{T_{n}(y)\left(\sum_{i=1}^{\infty}a_{i}T_{i}(y)\right)}{\pi\sqrt{1-y^{2}}}\mathrm{d}y\right)$
@@
$F(\mu)$ $=\log 2\left(\int_{-1}^{1}\frac{1}{\pi\sqrt{1-x^{2}}}\mathrm{d}x\right)^{2}$
$\quad+\sum_{n=1}^{\infty}\frac{2}{n}\left(\int_{-1}^{1}\frac{a_{n}T_{n}(x)T_{n}(x)}{\pi\sqrt{1-x^{2}}}\mathrm{d}x\right)\left(\int_{-1}^{1}\frac{a_{n}T_{n}(y)T_{n}(y)}{\pi\sqrt{1-y^{2}}}\mathrm{d}y\right)$
$=\log 2+\sum_{n=1}^{\infty}\frac{a_{n}^{2}}{2n}.$
@@
H.2 Additional Notes on the $\mathrm{O}(2m)\setminus\mathrm{SO}(2m)$ Case
@@
H.3 The Exact Form of Rate Function $I(r)$
@@
Like the Gross-Witten Ensemble (GrossWitten1980), there is a third-order phase transition near $r=1/2$.
Appendix I Layer-wise Homology
We compare our method against REEF (Representation Encoding Fingerprint) in revealing layer-wise homology (zhang2025reef) on the following two model pairs:
  • Llama-3.1-8B vs. Llama-3.2-1B (meta2024llama32): A pruned version where we expect a structured layer correspondence;
  • Llama-3.1-8B vs. Qwen3-8B-Base (yang2025qwen3technicalreport): Two independently developed models with no known lineage.
(a) MDIR vs. REEF on revealing layer-wise relasionship.
In Figure 10(a)(a), MDIR produces a diagonal-like pattern, indicating that each layer of Llama-3.2-1B aligns with a specific layer in Llama-3.1-8B (suggesting the 16 layers of Llama-3.2-1B are derived from layer 0, 1, 2, 3, 4, 5, 8, 11, 14, 17, 20, 23, 26, 29, 30 and 31 of Llama-3.1-8B, respectively). This reveals not only homology but also the exact strategy used during model pruning. This offers MDIR a very fine-grained level of interpretability which previous methods cannot provide. In contrast, REEF (Figure 10(a)(c)) yields uniformly high CKA similarity measures across many pairs of layers and failing to identify the exact correspondence. While it captures global similarity, it lacks the granularity to reveal which layers are actually aligned, thus less effective for model homology detection. When applied to unrelated models (Figure 10(a)(b,d)), MDIR correctly outputs a noise pattern, consistent with the absence of structural homology. REEF, however, continues to report high similarities (>0.9>0.9) between certain layer pairs. Our results demonstrate that MDIR is capable of interpretably reconstructing layer mappings between related model weights.